OpenStudy (anonymous):

Solve this equation for x. Round your answer to the nearest hundredth. 7=ln(x+3)Solve this equation for x. Round your answer to the nearest hundredth. 7=ln(x+3)

OpenStudy (unklerhaukus):

\[7=\ln(x+3)\]\[7=\log_e(x+3)\]\[e^7=e^{\log_e(x+3)}\]\[e^7=x+3\] \[x=?\]